How many different lists are there of length four consisting of numbers (1,2,3). Each number needs to appear at least once.
Answer #1

since each number appear at least once, there a number from (1,2,3)will appear twice and other only once
number of ways =N(select a number which appear twice from 3 numbers and then arrange the numbers ) =(3C2)*(4!/(2!*1!*1!)) =3*4*3 =36